Snookers Lumber can convert logs into either lumber or plywood. In a given day, the mill turns out three times as many units of plywood as lumber. It makes a profit of $30 on a unit of lumber and $40 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $15300??

How many units of lumber and how many units of plywood are needed to be sold to reach the profit??

Answer by jessica43(140)

To solve this problem, you are going to have to write 2 equations using what you know.
First, you know that the mill turns out 3 times as many units of plywood as lumber. This can be written as:
P = 3L (where P = units of plywood, L = units of lumber)
Second, you know that it makes a profit of $30 on a unit of lumber and $40 on a unit of plywood and you want the total profit to equal $15300. So:
30L + 40P = 15300
Now you can plug the first equation into the second equation and solve for L:
30L + 40P = 15300
30L + 40(3L) = 15300
30L + 120L = 15300
150L = 15300
L = 102
So 102 units of lumber must be produced and sold.
To find the number of plywood units, plug this into the first equation:
P = 3L
P = 3(102)
P = 306
So 306 units of plywood must be produced and sold.